Binomial probability:
A company takes a sample of 20 boxes of cereal from their production line. They believe that of the 20 boxes, four will have too little cereal in them and sixteen will have the correct amount. The company then randomly selects three of the boxes from the twenty and weighs them. Assuming the company is correct about the number of light boxes, find the probability that the three boxes selected will all have the correct amount.
If n=20 and r=3, what is the probability of success?

(I'm just having some trouble figuring out what exactly is Probability for Success in this math problem. Therefore, please help to figure out what the heck is P=? because than I will be able to figure out what q=? on my own, Thank you! Also, I really appreciate it too.)

Answer by stanbon(75887)

A company takes a sample of 20 boxes of cereal from their production line. They believe that of the 20 boxes, four will have too little cereal in them and sixteen will have the correct amount.
-----
p-hat = 16/20 = 4/5 (probability a box has the correct amount)
----------
The company then randomly selects three of the boxes from the twenty and weighs them. Assuming the company is correct about the number of light boxes, find the probability that the three boxes selected will all have the correct amount.
If n=20 and r=3, what is the probability of success?
-----
Binomial Problem with n=3 and p = 4/5
P(x=3) = (4/5)^3
---------------------
Cheers,
Stan H.
===========